Northwoods Backpackers is a retail catalog store in Vermont that specializes in outdoor clothing and camping...

  1. Northwoods Backpackers is a retail catalog store in Vermont that specializes in outdoor clothing and camping equipment. Phone orders are taken each day by a large pool of computer operators, some of whom are permanent and some temporary. A permanent operator can process an average of 76 orders per day, whereas a temporary operator can process an average of 53 orders per day. The company averages at least 600 orders per day. The store has 10 computer workstations. A permanent operator processes about 3 orders with errors each day, whereas a temporary operator averages 4.1 orders with errors daily. The store wants to limit errors to 24 per day. A permanent operator is paid $81 per day, including benefits, and a temporary operator is paid $50 per day. The company wants to know the number of permanent and temporary operators to hire to minimize costs.
Answer #1

Let the no. of permanent operators be “x” and temporary operators be “y”.

Total costs = 70x+50y. This is the objective function and has to be minimized.

Constraints:

1. 76x+53y>=600

2. x+y <= 10

3. 3x+4.5y<=24

4. x,y>=0 and should be integers

Using the solver function in excel the following solution is obtained:

x 8.00
y 0.00
Formula
Total costs 560.00 70x+50y
Constraints
608.00 >= 600.00 76x+53y>=600
8.00 <= 10.00 x+y <= 10
10.40 <= 24.00 3x+4.5y<=24

Thus minimized cost = $560 and the company will hire 8 permanent workers and 0 temporary workers. All constraints are satisfied. No. of orders processed = 608 (>600), no. of computer workstations used = 8(<10) and no. of errors = 10.4 (<24).

Solver image:

Clear 2 Show Detail ? Solver 21 AF Solver Parameters Connections Properties e Edit Links Connections 21 Sort Filter Reapply A